# Learning Streaming Video Representation via Multitask Training

### v1

@@ 4.1 Pre-training Data
We leverage various publicly available video datasets for training, encompassing annotations across various granularities. As shown in Table 1, these datasets are categorized based on the task, including action recognition (AR) [12, 36], video-text retrieval (VTR) [103, 14, 11, 2, 97, 84], temporal action localization (TAL) [11, 64, 121], temporal video grounding (TVG) [35, 49, 78, 48, 2, 69] , video instance segmentation (VIS) [108, 91, 57], referring video object segmentation (RVOS) [26, 80]. The total number of video-text pairs used for pre-training is $\sim$1M.
Task Pre-training Dataset Scale
Global-level
AR K400 [12], SSv2 [36] 400K
Temporal-level
TAL ActivityNet-1.3 [11], FineAction [64], HACS [121] 180K
Spatial-level
VIS YouTubeVIS-19 [108], LVVIS [91], COCO [57] 120K
RVOS MEVIS [26], Refer-YouTube-VOS [80] 36K
Total - $\sim$1M
Table 1: Pretraining datasets.
@@ 4.2 Implementation Details
We initialize the spatial attention layers using pre-trained weights from SigLIP [114], while the temporal attention layers are randomly initialized. To facilitate smooth integration of temporal information, we incorporate a learnable tanh gate (initialized to zero) in the temporal attention. This design ensures the model initially preserves SigLIP’s strong spatial encoding capabilities, and gradually learns temporal dependencies. At training time, we perform over-sampling to balance the data scale of different tasks. In all experiments, we use the base model (siglip-base-patch16-224). We train our model for 20 epochs, with global batch size 128, and a learning rate of $10^{-6}$. By default, we uniformly sample 16 frames from the video as input. After training StreamFormer jointly on multiple tasks, we transfer the model to downstream video tasks by fixing the backbone and training task-specific heads.
@@ 4.3 Comparison on Downstream Tasks
We evaluate StreamFormer on multiple downstream tasks, including online action detection, online video instance segmentation, and video question answering.
@@ 4.3.1 Online Action Detection
As shown in Table 2, our model achieves state-of-the-art results. Notably, our model demonstrates a significant improvement over the baseline SigLIP without using optical flow (58.9 vs. 68.4), and it even achieves comparable performance with early approaches that uses flow. Given that the extraction of optical flow is highly time-consuming [15, 105], our approach clearly offers more efficient online inference without relying on flow, while still ensuring a high accuracy in action detection.
@@
Method Res. Backbone COCO-init. Mask Head YTVIS19
@@ 4.3.2 Online Video Instance Segmentation
The goal of online video instance segmentation is simultaneous detection, segmentation, and tracking of instances in videos, in a frame-by-frame manner. With our causal temporal attention, we can evaluate our video backbone model like an image backbone. We choose CTVIS [111] as our baseline, and keep all hyper-parameters identical while changing the backbone into our StreamFormer. Due to the requirement for dense features in this task, we add ViT-Adapter [19, 118] to our backbone. Given the enormous computation cost for the traditional VIS backbone training from image instance segmentation [20, 21], we exclusively train video instance segmentation on video datasets only. We report the Average Precision (AP) and Average Recall on YoutubeVIS-2019 [108] benchmark.
@@ 4.3.3 Video Question Answering
Building upon the frameworks established by LLaVA [60] and LLaVA-Next-Video [119], we propose to integrate StreamFormer with large language models for video question answering. For a fair comparison, we replace the CLIP image encoder in LLaVA-Next with StreamFormer (frozen in all training stages), and keep the remaining architecture and training pipeline unchanged. We use the same image-text pairs in LLaVA-NeXT-Video, and sample 80K video-text pairs from LLaVA-Video-178K [120] . We provide our full training details in the supplementary materials.
Table 4 lists the comparison results on VideoMME [34] and MLVU [125]. StreamFormer exhibits superior performance among methods that use 224$\times$224 resolution, indicating that enhancing temporal motion and spatial detail perception in video backbone enables better VideoQA.
@@ 4.4 Ablation Study
Ablation study for pre-training tasks. Table 5 shows the model’s downstream performance when pre-trained on different task combinations. Our baseline is the original image encoder of SigLIP, i.e., without multitask training and temporal modeling. We draw the following conclusions: (1) The baseline achieves relatively good performance on downstream tasks, highlighting the strong image representation capability of SigLIP; (2) the incorporation of any form of video-language pre-training, either global, temporal, or spatial, yields a notable improvement over the baseline on all downstream tasks; (3) both global- and temporal-level tasks yield better results across various downstream tasks, while models trained solely on spatial-level tasks perform sub-optimally in online action detection and video question answering. This may be attributed to the relatively limited amount of data available for video segmentation, leading to insufficient understanding of verbs/actions; (4) by integrating tasks of varying granularities, the model achieves competitive downstream performance, proving the effectiveness of our multitask training approach.
@@
Effect of causal temporal modeling. We investigate the effect of causal temporal attention (Causal TA) by comparing with two baselines: (1) an image-based model that only employs spatial attention combined with LoRA, without any temporal attention (No TA); (2) a model utilizes bi-directional temporal attention (Bi-directional TA). We report the performance of the models without optical flow to verify the temporal effect of visual features. As observed from Table 6, temporal attention matters in online action detection, while bi-directional attention performs slightly better than causal attention. This is reasonable as each frame can access all frames in bi-directional attention.
@@ 5 Conclusions
In this paper, we present StreamFormer, a streaming video backbone that incorporates causal temporal attention and spatial low-rank adaptation into pre-trained SigLIP image encoder. To facilitate the training of StreamFormer, we devise a multitask visual-language alignment framework that unifies diverse video understanding tasks, enabling to learn global-level semantics, temporal-level dynamics and spatial-level perception simultaneously. We conduct extensive experiments on online action detection, online video instance segmentation and video question answering. StreamFormer exhibits promising online video understanding ability while preserving high efficiency, demonstrating its huge potential in future real-time applications.

### v2

@@ 4.1 Pre-training Data
We leverage various publicly available video datasets for training, encompassing annotations across various granularities.
The datasets are categorized based on the task, including action recognition (AR) [12, 38], video-text retrieval (VTR) [107, 14, 11, 2, 101, 88], temporal action localization (TAL) [11, 67, 126], temporal video grounding (TVG) [37, 52, 82, 51, 2, 72] , video object segmentation (VOS) [113, 95, 60], referring video object segmentation (RVOS) [26, 84]. The total number of video-text pairs used for pre-training is $\sim$1M.
@@ 4.2 Implementation Details
We initialize the spatial attention layers using pre-trained weights from SigLIP [119] with LoRA rank 32, while the temporal attention layers are randomly initialized. To facilitate smooth integration of temporal information, we incorporate a learnable tanh gate (initialized to zero) in the temporal attention. This design ensures the model initially preserves SigLIP’s strong spatial encoding capabilities, and gradually learns temporal dependencies. At training time, we perform over-sampling to balance the data scale of different tasks. In all experiments, we use the base model (siglip-base-patch16-224). We train our model for 20 epochs, with global batch size 128, and a learning rate of $1e^{-6}$. By default, we uniformly sample 16 frames from the video as input. After training StreamFormer jointly on multiple tasks, we transfer the model to downstream video tasks by fixing the backbone and training task-specific heads.
@@ 4.3 Comparison on Downstream Tasks
We freeze StreamFormer and evaluate it on multiple downstream tasks, including online action detection, online video instance segmentation, and video question answering.
@@ 4.3.1 Online Action Detection
As shown in Table 1, our model achieves state-of-the-art results, demonstrating a significant improvement over the baseline SigLIP without using optical flow (58.9 vs. 68.4), and even achieves comparable performance with early approaches that use flow. Given the time-consuming nature of extracting optical flow [16, 110], our approach clearly offers more efficient online inference without relying on flow, while still ensuring a high accuracy in action detection.
@@
Method Res. Backbone COCO-init. YTVIS19
@@ 4.3.2 Online Video Instance Segmentation
The goal of online video instance segmentation is the simultaneous detection, segmentation, and tracking of instances in videos, in a frame-by-frame manner. With our causal temporal attention, we can evaluate our video backbone model like an image backbone. We choose CTVIS [116] as our baseline, and keep all hyperparameters identical while changing the backbone into frozen StreamFormer.
Due to the requirement for dense features in this task, we add ViT-Adapter [20, 123] to our backbone. Note that we do not initialize model parameters with pre-training on COCO [60], and instead directly train YoutubeVIS-2019 [113] with COCO joint training.
We report the Average Precision (AP) and Average Recall on YoutubeVIS-2019 benchmark.
@@ 4.3.3 Video Question Answering
Building upon the frameworks established by LLaVA [63] and LLaVA-Next-Video [124], we propose to integrate StreamFormer with large language models for video question answering. For a fair comparison, we replace the CLIP in LLaVA-Next (Vicuna-7b) with StreamFormer (frozen in all training stages), and keep the remaining architecture and training pipeline unchanged. We use the same image-text pairs in LLaVA-NeXT, and sample 80K video-text pairs from LLaVA-Video-178K [125] . We provide our full training details in the supplementary materials.
Table 3 presents the comparison results on VideoMME [36] and MLVU [132]. StreamFormer exhibits superior performance among methods that use 224$\times$224 resolution, indicating that enhancing temporal motion and spatial detail perception in video backbone enables better VideoQA.
@@ 4.4 Ablation Study
Ablation study for pre-training tasks. Table 4 shows the model’s downstream performance when pre-trained on different task combinations. Our baseline is the original image encoder of SigLIP, i.e., without multitask training and temporal modeling. We draw the following conclusions: (1) the baseline achieves relatively good performance on downstream tasks, highlighting the strong image representation capability of SigLIP; (2) the incorporation of any form of video-language pre-training, either global, temporal, or spatial, yields a notable improvement over the baseline on all downstream tasks; (3) both global- and temporal-level tasks yield better results across various downstream tasks, while models trained solely on spatial-level tasks perform sub-optimally in online action detection and video question answering. This may be attributed to the relatively limited amount of data available for video segmentation, leading to an insufficient understanding of verbs/actions; (4) by integrating tasks of varying granularities, the model achieves competitive downstream performance, proving the effectiveness of our multitask training approach.
@@
Effect of temporal causal modeling. We investigate the effect of temporal causal attention (Causal TA) by comparing it with two baselines: (1) an image-based model that only employs spatial attention combined with LoRA, without any temporal attention (No TA); (2) a model utilizes bi-directional temporal attention (Bi-directional TA). We report the performance of the models without optical flow to verify the temporal effect of visual features. As observed from Table 5, temporal attention matters in online action detection, while bi-directional attention performs slightly better than causal attention. This is reasonable as each frame can access all frames in bi-directional attention.
@@
[…]
We randomly select 1M video-text pairs from WebVid-10M, which is comparable to the scale of our pre-training data. As shown in Table 6, the models trained on WebVid-1M exhibit relatively low performance, possibly due to insufficient pre-training data for video-text contrastive learning. In comparison, our approach outperforms WebVid-1M model using only 0.1M pre-training data, significantly reducing the computational cost for training. Our method effectively leverages the existing high-quality annotated video datasets to enhance the backbone’s capabilities, which suggests there is certainly a trade-off between computation and annotation cost.
@@ 5 Conclusions
In this paper, we present StreamFormer, a streaming video backbone that incorporates causal temporal attention and spatial low-rank adaptation into a pre-trained SigLIP image encoder. To facilitate the training of StreamFormer, we devise a multitask visual-language alignment framework that unifies diverse video understanding tasks, enabling it to learn global-level semantics, temporal-level dynamics and spatial-level perception simultaneously. We conduct extensive experiments on online action detection, online video instance segmentation and video question answering. StreamFormer exhibits promising online video understanding ability while preserving high efficiency, demonstrating its huge potential in future real-time applications.
Acknowledgement
Weidi would like to acknowledge the funding from Scientific Research Innovation Capability Support Project for Young Faculty (ZY-GXQNJSKYCXNLZCXM-I22).
@@ Acknowledgement
1 Pre-training Datasets
We list StreamFormer’s pre-training datasets in Table  7. Note that for our Video Object Segmentation datasets, we utilize the VIS datasets and VIS-style produced COCO-pseudo videos for their relatively well correspondence within the objects inside the videos or pseudo videos.
Task Pre-training Dataset Scale
Global-level
AR K400 [12], SSv2 [38] 400K
Temporal-level
TAL ActivityNet-1.3 [11], FineAction [67], HACS [126] 180K
Spatial-level
VOS YouTubeVIS-19 [113], LVVIS [95], COCO [60] 120K
RVOS MEVIS [26], Refer-YouTube-VOS [84] 36K
Total - $\sim$1M
Table 7: Pretraining datasets. The abbreviations are defined as follows: “AR” for action recognition, “VTR” for video-text retrieval, “TAL” for temporal action localization, “TVG” for temporal video grounding, “VOS” for video object segmentation, and “RVOS” for referring video object segmentation.
@@ 2.3 Video Question Answering
3 More Results
We add more results of StreamFormer in this section.
3.1 Video Action Recognition
In Table 10, we add more comparisons with recent causal and autoregressive models, including TRecViT [32] and Toto [81] on Video Action Recognition tasks of K400 and SSv2. While StreamFormer achieves stronger results on K400 and is competitive on SSv2, we emphasize that StreamFormer’s key advantage is multitask learning across different spatio-temporal granularities—from short-term action recognition to fine-grained, frame-level tasks like OAD and Online VIS.
3.2 Downstream Evaluations
In Table 10, we continue to supplement more results with TVSeries [24] (w/o flow) for Online Action Detection and OVIS [76] (224p and w/o COCO pre-training) for Online Video Instance Segmentation.In Table 11, we detail our results in Real-Time Visual Understanding task of StreamingBench [31]. Note that our LLM is Vicuna-7B-v1.5.
3.3 Qualitative Results
We also add qualitative results of StreamFormer’s Online Video Instance Segmentation in Figure  4. Frames are sampled from the validation set of Youtube-VIS 2019.
Backbone K400 SSv2
Toto 64.7 -
TRecViT 78.4 66.8
StreamFormer 82.2 66.5
Table 9: Comparison to other causal backbones.
Backbone TVSeries OVIS
mAP AP
SigLIP 84.8 18.2
StreamFormer 88.1 20.8
Table 10: Comparison on additional downstream datasets.
Figure 4: Qualitative results of Online Video Instance Segmentation. Zoom in for a better view.
  Model Frames StreamingBench Real-Time Visual Understanding OP CR CS ATP EU TR PR SU ACP CT All   InternVL-V2 16 68.12 60.94 69.40 77.12 67.70 62.93 59.26 53.25 54.96 56.48 63.72 Kangaroo 64 71.12 84.38 70.66 73.20 67.08 61.68 56.48 55.69 62.04 38.86 64.60 LongVA 128 70.03 63.28 61.20 70.92 62.73 59.50 61.11 53.66 54.67 34.72 59.96 VILA-1.5 14 53.68 49.22 70.98 56.86 53.42 53.89 54.63 48.78 50.14 17.62 52.32 Video-CCAM 96 56.40 57.81 65.30 62.75 64.60 51.40 42.59 47.97 49.58 31.61 53.96 Video-LLaMA2 32 55.86 55.47 57.41 58.17 52.80 43.61 39.81 42.68 45.61 35.23 49.52   LLaVA-Next (StreamFormer) 16 69.16 66.67 63.10 57.40 70.2 50.00 67.86 52.8 66.67 12.5 58.24
Table 11: Performance comparison on StreamingBench Real-Time Visual Understanding tasks. The abbreviations are defined as follow: “OP” for Object Perception, “CR” for Causal Reasoning, “CS” for Clips Summarization, “ATP” for Attribute Perception, “EU” for Event Understanding, “TR” for Text-Rich Understanding, “PR” for Prospective Reasoning, “SU” for Spatial Understanding, “ACP” for Action Perception and “CT” for Counting.